Solve the following equations:
2q + 6 =12 …...........

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

First we have to subtract 6 to the both sides of the equation,
Then, we get,
= 2q + 6 - 6 = 12 - 6
= 2q = 6
Now,
We have to divide both sides of the equation by 2,
Then, we get,
= 2q/2 = 6/2
= q = 3

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows :
2q + 6 = 12
2q = 12 - 6
2q = 6
q= 6/2
q = 3
 
Hope it helps!
Thank you ✌️